# Always-Sparse Training by Growing Connections with Guided Stochastic Exploration

### v1

@@ Appendix I Extended training
The following extended training results are for CIFAR-100 at 98% sparsity with 1x, 1.5x, and 2x the number of training epochs. The learning rate drops and the final prune-grow step are also scaled by the same amount, all other hyperparameters stay the same. It is clear that all models benefit significantly from extended training. The ViT model in particular gains 7% in accuracy by training for twice as long.
| Epochs | 200 (1x) | 300 (1.5x) | 400 (2x) |
@@
Appendix J Model scaling training progress
In Figure 7, we provide plots for the test and training accuracy during training for our method on CIFAR-100 while increasing the width of the model by 1x, 2x, 3x, and 4x. The number of active connections is kept constant between all experiments, see Section 4.5 for details. In the CNN models, ResNet-56 and VGG-16, the increase in width causes the models to train significantly faster and obtain higher accuracy. For the vision transformer (ViT) model, the trend is reversed, wider models train slower and obtain lower accuracy.
Appendix K Scaling dense ViT
To further investigate the behavior of ViT, which decreases in accuracy when using larger models with the same number of activate parameters, as discussed in Section 4.5 and Appendix J. In Table 4, we report the accuracy of ViT for increasing model width but kept dense, thus increasing the active connections count. We see the accuracy quickly plateaus at 62.2%, gaining only 2% accuracy while even the sparse CNNs gained 6.5% on average at 4 times wider. This indicates that the decreasing accuracy of ViT reported in Section 4.5 is not due to our sparse training method as even the dense model does not benefit much from increased scale.
| Model width | 1x | 2x | 3x | 4x |
| Accuracy | 60.3±0.6 | 62.5±0.5 | 62.2±0.4 | 62.2±0.1 |
Appendix L Sparse matrix multiply benchmark
In Figure 8, we show the execution time of the sparse times dense matrix-matrix multiply on two execution devices: Intel Xeon Gold 6148 CPU at 2.40GHz and NVIDIA Tesla V100 16GB. The sparse matrix is a square matrix of size units times units, representing an unstructured sparse random weight matrix. The dense matrix is of size units times 128 and represents a batch of layer inputs. We compare three storage formats, namely: dense, where all the zeros are explicitly represented; Coordinate format (COO), which represents only the nonzero elements and their coordinates; and Compressed Sparse Row (CSR), which is similar to COO but compresses the row indices. Each data point in Figure 8 is the average of 10 measurements.
The results show that the CSR format generally outperforms the COO format. Moreover, the execution time of the CSR format improves upon the dense execution time starting around 90% sparsity. At 99% sparsity, the CSR format is up to an order of magnitude faster than the dense format. These results are similar between the two execution devices, although the NVIDIA Tesla V100 16GB is up to two orders of magnitude faster than the Intel Xeon Gold 6148 CPU at 2.40GHz.
Appendix M Layer-wise sparsity
In Figure 9, we show the sparsity of each layer, in terms of compression, obtained by each sparsification methods at the end of training. In the plots the CIFAR-10 and CIFAR-100 results are averaged. These plots can be used to gain insight into how each sparsification method distributes the active connections over the models.

### v2

@@ IV-E Extended training
We continued our experimentation by looking at the effect of extended training. The following extended training results are for CIFAR-100 at 98% sparsity with 1x, 1.5x, and 2x the number of training epochs. The learning rate drops and the final prune-grow step are also scaled by the same amount, all other hyperparameters stay the same. We observe that ViT and CNN models benefit considerably from extended training, with an average increase in accuracy of 5.3% for training twice as long. The ViT model in particular gains 7% in accuracy by training twice as long.
| Epochs | 200 | 300 | 400 |
@@
IV-F Floating-point operations
Lastly, to determine the training efficiency of GSE, we compare its FLOPs with those of RigL in Figure 4. FLOPs are determined for ResNet-50 on ImageNet using the method described by Evci et al. [35], with . The results show that the benefit of GSE increases rapidly at high sparsity levels. At 99% sparsity, GSE uses 13.3% fewer FLOPs than RigL, up from 2.4% at 95% sparsity.
V Conclusion
We presented an always-sparse training algorithm using guided stochastic exploration of the connections which has a linear time complexity with respect to the model width and improves upon the accuracy of previous sparse training methods. This is achieved by only evaluating gradients for a subset of randomly sampled connections when changing the connections. The three distributions for sampling the subset—uniform, an upper bound of the gradient magnitude, and an estimate of the gradient magnitude—all showed similar trends. However, uniform consistently achieves among the highest accuracy and is also the most efficient. We therefore conclude that the uniform distribution is the most appropriate for sparse training. Interestingly, we showed that training larger and sparser CNN models results in higher accuracy for the same number of active connections. Moreover, it was observed that our method greatly benefits from extended training. Lastly, the number of floating-point operations of our method decrease considerably faster than those for RigL as the sparsity increases. This observation underscores the efficiency and scalability of our method, which is promising for training increasingly larger and sparser ANNs.